OpenStudy (anonymous):

how do you graph and solve this quadratic function? xsquared-x+y+6=0

OpenStudy (mertsj):

\[x^2-x+y+6=0\] \[y+6=-x^2+x\] \[y+6=-1(x^2-x)\]

OpenStudy (anonymous):

isn't it supposed to be in standard form?

OpenStudy (mertsj):

\[y+6-\frac{1}{4}=-1(x^2-x+\frac{1}{4})\] \[y+\frac{23}{4}=-(x-\frac{1}{2})^2\] \[y=-(x-\frac{1}{2})^2-\frac{23}{4}\]